OpenStudy (anonymous):

(3^(sqrt(2)))^(-sqrt(2)) Can someone help simplify? I know that the answer is (1/9), but I'm having trouble getting there.

OpenStudy (anonymous):

\[(x ^{ n}) ^ {m}= x ^{n \times m}\] \[x^{- z} = \frac{1}{x^z}\]

OpenStudy (anonymous):

So -- \[3^{\sqrt{2}*-\sqrt{2}} = 3^{-2} = 1/9\]

OpenStudy (anonymous):

thanks